OpenStudy (anonymous):

Consider the following pair of equations: y = x + 4 y = –2x – 2 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (kidrah69):

use substitution

OpenStudy (kidrah69):

(x+4)=-2x-2

OpenStudy (kidrah69):

combine the like terms like we did in the other problem

OpenStudy (aaronandyson):

y = x + 4 - eqn 1 y = –2x – 2 - eqn 2 Put y = x + 4 in eqn 2 x + 4 = -2x - 2 Solve for x

OpenStudy (aaronandyson):

Hint: Combine the like terms.

OpenStudy (kidrah69):

thats exactly what I said

OpenStudy (aaronandyson):

x + 4 = -2x - 2 Isolate x x + 2x = -2 - 4

OpenStudy (aaronandyson):

@ladybugg98

OpenStudy (aaronandyson):

Are you there @ladybugg98 ?